A survey collected data on annual credit card charges in seven different categories of expenditures: transportation, groceries, dining out, household expenses, home furnishings, apparel, and entertainment. Using data from a sample of 42 credit card accounts, assume that each account was used to identify the annual credit card charges for groceries (population 1) and the annual credit card charges for dining out (population 2). Using the difference data, with population 1 − population 2, the sample mean difference was

d = $850,

and the sample standard deviation was

sd = $1,122.

(a)

Formulate the null and alternative hypotheses to test for no difference between the population mean credit card charges for groceries and the population mean credit card charges for dining out.

H0: μd ≠ 0

Ha: μd = 0

H0: μd = 0

Ha: μd ≠ 0

    

H0: μd ≤ 0

Ha: μd > 0

H0: μd ≥ 0

Ha: μd < 0

H0: μd < 0

Ha: μd = 0

(b)

Calculate the test statistic. (Round your answer to three decimal places.)

What is the p-value? (Round your answer to four decimal places.)

Can you conclude that the population means differ? Use a 0.05 level of significance.

The p ≤ 0.05, therefore we can conclude that there is a difference between the annual population mean expenditures for groceries and for dining out.The p > 0.05, therefore we can conclude that there is a difference between the annual population mean expenditures for groceries and for dining out.    The p ≤ 0.05, therefore we cannot conclude that there is a difference between the annual population mean expenditures for groceries and for dining out.The p > 0.05, therefore we cannot conclude that there is a difference between the annual population mean expenditures for groceries and for dining out.

(c)

What is the point estimate (in dollars) of the difference between the population means?

$

What is the 95% confidence interval estimate (in dollars) of the difference between the population means? (Round your answers to the nearest dollar.)

$  to $

Which category, groceries or dining out, has a higher population mean annual credit card charge?

The 95% confidence interval  ---Select--- is completely below is completely above contains zero. This suggests that the category with higher mean annual expenditure is
Answer #1

Ans:

a)

H0: μd = 0

Ha: μd ≠ 0

b)

Test statistic

t=(850-0)/(1122/SQRT(42))

t=4.910

df=42-1=41

p-value=tdist(4.910,41,2)=0.0000

c)As,p-value<0.05,we reject H0.

The p ≤ 0.05, therefore we can conclude that there is a difference between the annual population mean expenditures for groceries and for dining out.

c)Point estimate for difference=850

margin of error=2.0195*1122/SQRT(42)=350

95% confidence interval for difference

=850+/-350

=(500, 1200)

The 95% confidence interval is completely above zero. This suggests that the category with higher mean annual expenditure is groceries.